OpenStudy (anonymous):

Find the x and y intercepts 1.) 4y-8=2x

hartnn (hartnn):

To get x-intercept put y=0 in the equation to get y-intercept put x=0 in the equation

OpenStudy (anonymous):

can you help me do it and explain each step

hartnn (hartnn):

4y-8=2x FOR x-INTERCEPT, PUT y=0 in that, what u get ?

OpenStudy (anonymous):

 that is what i got

OpenStudy (tkhunny):

1.) 4y-8=2x Add 8 4y = 8 + 2x Subtract 2x 4y - 2x = 8 Divide by 8 \[\frac{y}{2} + \frac{x}{-4} = 1\] This is called the "Intercept Form". The answers are staring at you.

hartnn (hartnn):

-8=2x so x= -8/2 = ?

OpenStudy (tkhunny):

You have it. Now the other?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

hartnn

hartnn (hartnn):

that is correct.

OpenStudy (tkhunny):

Okay, now start with x = 0 and find the y-intercept.

hartnn (hartnn):

4y-8=0 4y=8 y=?

OpenStudy (anonymous):

2

OpenStudy (tkhunny):

Now go take a look at that Intercept Form and see if anything looks interesting - particularly the denominators.

hartnn (hartnn):

thats correct. any more doubts ?